OpenStudy (anonymous):

If I am given a triangle with all 3 angles but only 1 side length how do I find the other sides?

OpenStudy (anonymous):

trigonometric identities?

OpenStudy (anonymous):

you can draw height to the triangle then use basic trigonometry

OpenStudy (phi):

Law of Sines

OpenStudy (anonymous):

How do I use law of sines?

OpenStudy (phi):

write the ratio of sin(A)/opposite side = sin(B)/opposite side if we label the side opposite angle A as a, (i.e. use lower case letters for the sides) use \[ \frac{\sin A}{a} = \frac{\sin B}{b} \]

OpenStudy (anonymous):

Idk what to do after this:\[\frac{ \sin(83) }{ a }=\frac{ \sin(45) }{ b }=\frac{ \sin(52) }{ 12 }\]

OpenStudy (phi):

just do one pair at a time \[ \frac{ \sin(83) }{ a }=\frac{ \sin(52) }{ 12 } \] one way is first cross multiply. can you do that ?

OpenStudy (anonymous):

I got a = 15.08?

OpenStudy (phi):

you should get a= 12*sin(83)/sin(52)= 15.11

OpenStudy (anonymous):

Okay yea I redid it and got that

OpenStudy (anonymous):

I got 15.08 by rounding

OpenStudy (phi):

ok, now do the other pair \[ \frac{ \sin(45) }{ b }=\frac{ \sin(52) }{ 12 } \]

OpenStudy (anonymous):

a=10.77

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay thank you!